OpenStudy (kkutie7):

Runner run in a straight line with the average velocity of 5m/s for 5min and then with an average velocity of 4m/s 4 minutes. Total displacement is 2460m. What is the average velocity during this time?

OpenStudy (abhisar):

Average velocity = Total Displacement/Total time taken => 2460/9*60 = 4.55 m/s

OpenStudy (kkutie7):

suppose the goal is to run 5 km in 15 minutes. What average velocity must she have over the remaining portion of her run?

OpenStudy (kkutie7):

\[5km*\frac{1000m}{1km}=5000m\] \[15min*\frac{60s}{1min}=900s\]

OpenStudy (kkutie7):

@Abhisar

OpenStudy (abhisar):

In continuation to the first question ?

OpenStudy (kkutie7):

I guess. I have the answer if that helps. I just need to know how to get there.

OpenStudy (abhisar):

ok, what's the answer ?

OpenStudy (kkutie7):

7.06m/s

OpenStudy (kkutie7):

I was thinking that maybe 5000m would be the new displacement. Or to subtract 5000m-2460m=2540m would be the new displacement... I'm not sure where I'm going with this

OpenStudy (abhisar):

Since, runner has to run a total of 5000m and she/he has already covered 2540m. We will consider a displacement of 5000-2540m. Also he/she has to complete the whole 5000m in 15 mins. But he/she has already taken 5+4=9 mins. So he/she will have to complete 2550 m in 15-9=6 mins. Velocity = Displacement/Time => 2550/6*60 = 7.08 m/s

OpenStudy (kkutie7):

where did you get 2250?

OpenStudy (abhisar):

Sorry it's 2460

OpenStudy (kkutie7):

oh ok =) thanks!

OpenStudy (abhisar):

No ! lol It's 2540, 5000-2460=2540

OpenStudy (kkutie7):

oh shoot haha

OpenStudy (abhisar):

So everything's clear now ?

OpenStudy (kkutie7):

very thank you =D

OpenStudy (abhisar):

you're most welcome :)

OpenStudy (abhisar):

Just to clear it out, the final step will be little tweaked. Velocity = Displacement/Time => 2540/6*60 = 7.06 m/s

OpenStudy (kkutie7):

yupp got it